# AstroMLab 4: Benchmark-Topping Performance in Astronomy Q&A with a 70B-Parameter Domain-Specialized Reasoning Model

### v1

@@ Abstract
General-purpose large language models (LLMs), despite their broad capabilities, often struggle with specialized domain knowledge, a limitation particularly pronounced in more accessible, lower-parameter versions. This gap hinders their deployment as effective agents in demanding fields such as astronomy. Building on our prior work with AstroSage-Llama-3.1-8B, this study introduces AstroSage-Llama-3.1-70B, a significantly larger and more advanced domain-specialized natural-language AI assistant. It is designed for research and education across astronomy, astrophysics, space science, astroparticle physics, cosmology, and astronomical instrumentation. Developed from the Meta-Llama-3.1-70B foundation, AstroSage-Llama-3.1-70B underwent extensive continued pre-training (CPT) on a vast corpus of astronomical literature, followed by supervised fine-tuning (SFT) and model merging. Beyond its 70-billion parameter scale, this model incorporates refined datasets, judiciously chosen learning hyperparameters, and improved training procedures, achieving state-of-the-art performance on complex astronomical tasks. Notably, we integrated reasoning chains into the SFT dataset, enabling AstroSage-Llama-3.1-70B to either answer the user query immediately, or first emit a human-readable thought process. Evaluated on the AstroMLab-1 benchmark [1]—comprising 4,425 questions from literature withheld during training—AstroSage-Llama-3.1-70B achieves state-of-the-art performance. It surpasses all other tested open-weight and proprietary models, including leading systems like o3, Gemini-2.5-Pro, Claude-3.7-Sonnet, Deepseek-R1, and Qwen-3-235B, even those with API costs two orders of magnitude higher. This work demonstrates that domain specialization, when applied to large-scale models, can enable them to outperform generalist counterparts in specialized knowledge areas like astronomy, thereby advancing the frontier of AI capabilities in the field.
@@
I Introduction
Astronomy and its related fields demand sophisticated tools that can process vast amounts of specialized knowledge. Large Language Models (LLMs) have emerged as promising assistants for this domain, offering capabilities as research collaborators, educational resources, and knowledge repositories. Domain-specialized models demonstrate particular cost-effectiveness in such contexts, as their parameters can be optimized for specific knowledge domains rather than distributed across the breadth of general internet content. Our previous work, AstroSage-Llama-3.1-8B [2] (hereafter TdH25), established that a relatively modest 8-billion parameter LLM, when extensively trained on astronomical content, could match or exceed the performance of much larger general-purpose models on astronomical knowledge tasks [1]. This finding highlighted the potential of domain specialization for creating efficient, high-performing AI assistants.
@@ I Introduction
This paper is structured as follows: Section II details the multi-stage training pipeline for AstroSage-Llama-3.1-70B, including improvements to our CPT and SFT corpora and procedures, and the model merging strategy for the final release. Section III discusses potential applications. Section IV evaluates AstroSage-Llama-3.1-70B against the updated Ting et al. [1] benchmark, comparing it with other recent proprietary and open-weight models.
II Model Architecture and Training Methodology
AstroSage-Llama-3.1-70B is derived from the Meta-Llama-3.1-70B architecture [3]. This base model was selected for consistency with TdH25, which chose Meta-Llama-3.1-8B for its state-of-the-art general capabilities and permissive licensing. The tokenizer from Meta-Llama-3.1-70B-Instruct was used without modification. Following the methodology of TdH25, our development process comprised three main stages: continued pre-training (CPT), supervised fine-tuning (SFT), and model merging.
II.1 Continued Pre-training (CPT)
The objective of CPT is to imbue the base model with extensive domain-specific knowledge from astronomical literature. The CPT dataset for AstroSage-Llama-3.1-70B builds upon the comprehensive corpus developed in TdH25, which included approximately 250,000 arXiv preprints (astro-ph, gr-qc from 2007-2024), nearly 30,000 Wikipedia articles related to astronomy, and internet-available textbooks. The knowledge cutoff for the astro-ph papers included in the CPT dataset remains January 2024.
@@ II.1 Continued Pre-training (CPT)
  • We applied ftfy [4] for consistent Unicode text normalization and some rule-based repetition removal to correct OCR failures, supplementing the perplexity-based cleaning methods detailed in TdH25.
  • To preserve general language understanding and mitigate catastrophic forgetting due to specialization, we incorporated a random selection of samples from the FineWeb dataset [5] into each training epoch. This addition of previous pretraining tokens during CPT is sometimes known as “replay”. Notably, the specific FineWeb samples were varied for each epoch, ensuring diverse exposure to general web text.
The CPT phase was conducted on the Oak Ridge Leadership Computing Facility (OLCF) Frontier supercomputer utilizing AMD Instinct MI250X GPUs. The Meta-Llama-3.1-70B base model underwent 2.5 epochs of CPT, an extensive training process that utilized approximately 168,000 GPU-hours. Several learning hyperparameters were modified compared to TdH25. Firstly, the learning rate was scaled relative to that in Dubey et al. [3] to maintain the same learning rate per token. Secondly, the warm-up period was extended to 0.15 epochs. Finally, we increased the weight decay regularization term to 0.1, aligning with Dubey et al. [3].
II.2 Supervised Fine-tuning (SFT)
@@
Figure 1: Composition of the AstroSage-Llama-3.1-70B SFT training dataset. The combination of reasoning-focused datasets (41.8%) with domain-specific astronomy Q&A (30.8%) reflects our strategy to develop a model that combines analytical thinking with specialized knowledge. General instruction-following data (OpenHermes 2.5) helps maintain versatility while preserving domain expertise. File sizes represent uncompressed UTF-8 encoded text including system and user prompts.
Figure 1 illustrates the composition of the SFT dataset. Its largest component is NVIDIA’s Llama-Nemotron-Post-Training-Dataset [6], which was used to train the Llama-3_1-Nemotron series. This series consistently demonstrates excellent performance on the LMArena [7], suggesting it is a strong dataset for eliciting reasoning and aligning with human preferences. The reasoning components cover science, code, mathematics, and general chat, providing a foundation for analytical thinking across different domains.
@@
To enhance domain expertise, we incorporated the custom domain-specific Q&A datasets from both TdH25 and de Haan [8], which together comprise approximately 30% of the training data.
@@
The model was fine-tuned on this SFT dataset for 0.6 epochs. This phase consumed approximately 13,000 GPU-hours on the same MI250X GPU architecture. Hyperparameters mirrored those of the CPT stage, with the exception of weight decay, which was removed.
II.3 Training Implementation
@@
The CPT and SFT stages were conducted on the Oak Ridge Leadership Computing Facility (OLCF) Frontier supercomputer using AMD Instinct MI250X GPUs. Our implementation employed the GPT-NeoX framework [9], which we adapted for compatibility with the Llama-3.1 architecture. Training was distributed across 2,048 Graphics Compute Dies (GCDs) using a multi-dimensional parallelism strategy: tensor parallelism 8, pipeline parallelism 8, and data parallelism 32. As GPT-NeoX does not currently support DeepSpeed ZeRO stage 2/3 with pipeline parallelism, we used ZeRO stage 1 [10] with activation checkpointing enabled. This configuration achieved a computational throughput of approximately 50 TFLOPS/s per GCD, consistent with performance metrics reported by Dash et al. [11].
@@
Figure 2: Training dynamics for continued pre-training (CPT) and supervised fine-tuning (SFT). The top panel shows the training loss trajectory across 2.5 epochs of CPT followed by 0.6 epochs of SFT. Despite an early significant spike during CPT Epoch 1, the loss steadily decreases throughout training with minimal discontinuity at epoch boundaries, indicating effective learning without overfitting. The SFT phase exhibits a rapid initial decrease in loss, eventually reaching below 0.6. The bottom panel reveals the learning rate schedule, including the initial warm-up period, several manual adjustments during early CPT (visible as step changes), the planned cosine decay during the final partial CPT epoch, and a separate warm-up and decay cycle for the SFT phase. Early termination of both phases reflects the computational resource constraints.
@@
Figure 2 illustrates the training dynamics during both CPT and SFT phases. Several noteworthy patterns emerge from these curves. The CPT loss exhibits a consistent downward trajectory without discontinuities at epoch boundaries—a positive indicator that the model was still learning effectively from the dataset without overfitting. This suggests that additional training epochs would likely have yielded further performance improvements, had computational resources permitted.
Resource constraints limited our training to 2.5 epochs for CPT and 0.6 epochs for SFT, resulting in an incomplete final annealing stage and preventing full convergence of the SFT phase. Despite these limitations, AstroSage-Llama-3.1-70B demonstrates superior performance in astronomy Q&A benchmarks compared to other leading models. With additional compute resources, we anticipate further performance gains as the training could be extended to leverage the full potential of our dataset.
@@
The SFT training loss initially exceeds the final CPT loss but rapidly decreases, ultimately dropping below 0.6. This indicates that, on average, the model assigns more than half its probability ($P\right>=\exp(-\mathrm{loss})$) to the correct token—likely reflecting the more predictable nature of assistant responses compared to the CPT corpus. The model obtained after SFT is named AstroSage-70B-SFT.
II.4 Model Merging
@@
To create the final, publicly released AstroSage-Llama-3.1-70B, we employed model merging (also known as parameter averaging) using the mergekit library [12]. This technique allows us to combine the strengths of our specialized SFT model with the robust instruction-following capabilities of other popular fine-tuned Llama-3.1-70B variants.
@@
The final mixture for AstroSage-Llama-3.1-70B was created using the DARE-TIES method [13] with parameters rescale: true and lambda: 1.2. The AstroSage-70B-CPT model (our domain-specialized, pre-instruction-tuning model) was designated as the “base model” for the merge operation. The components and their contributions are:
  • 70% AstroSage-70B-SFT (density 0.7)
  • 15% Llama-3.1-Nemotron-70B-Instruct (density 0.5)
  • 7.5% Meta-Llama-3.3-70B-Instruct (density 0.5)
  • 7.5% Meta-Llama-3.1-70B-Instruct (density 0.5)
This model mixture and its associated hyperparameters were determined through an iterative trial-and-error process, aiming to maximize performance on an internal set of astronomy-focused questions, distinct from the benchmark we use in Section IV. Consequently, the final model inherits the astronomical expertise from AstroSage-70B-SFT while incorporating the refined general instruction-following and conversational polish from the other constituent models.
III Features and Reasoning Capability
@@
A notable feature of AstroSage-Llama-3.1-70B is its explicit reasoning capability. This aligns with recent advances in the broader LLM field, where explicit reasoning has emerged as a critical development for enhancing model performance on complex tasks. The integration of reasoning mechanisms has become increasingly common in state-of-the-art models, including OpenAI’s o1 through o4 series, Anthropic’s Claude models with “thinking mode,” DeepSeek-R1, and others. These developments demonstrate that exposing and structuring the internal reasoning process allows models to tackle complex problems more systematically, resulting in improved accuracy and reliability.
@@
Building on these industry-wide insights, AstroSage-Llama-3.1-70B implements explicit reasoning that can be activated at inference time by setting the system prompt to detailed thinking on and prefilling the assistant completion with <think>. When enabled, the model generates a step-by-step reasoning process before providing the final answer to the user’s query. This is particularly beneficial for complex astronomical problems requiring multi-step analysis. As the reasoning tokens are enclosed within <think> and </think> tags, they can easily be hidden from the end-user if desired.
IV Evaluation

### v2

@@ Abstract
General-purpose large language models (LLMs), despite their broad capabilities, often struggle with specialized domain knowledge. This gap hinders their deployment as reliable research agents in demanding fields such as astronomy. Building on our prior work with AstroSage-Llama-3.1-8B, this study introduces AstroSage-Llama-3.1-70B, a 70-billion parameter domain-specialized natural-language AI assistant. It is designed for research and education across astronomy, astrophysics, space science, astroparticle physics, cosmology, and astronomical instrumentation. Developed from the Meta-Llama-3.1-70B foundation, AstroSage-Llama-3.1-70B underwent extensive continued pre-training (CPT) on a vast corpus of astronomical literature, followed by supervised fine-tuning (SFT) and model merging. We integrated reasoning chains into the SFT dataset, enabling AstroSage-Llama-3.1-70B to either answer the user query immediately, or first emit a human-readable thought process. Evaluated on a validated subset of 3,846 questions from the AstroMLab-1 benchmark (tingAstroMLab1Who2024)—derived from literature withheld during training—AstroSage-Llama-3.1-70B achieves top-tier performance (89.0%), matching GPT-5.2, Claude-4.5-Opus, and Gemini-3-Pro while being more cost-efficient. This work demonstrates that domain specialization, when applied to large-scale models, can enable them to outperform generalist counterparts in specialized knowledge areas like astronomy, thereby advancing the frontier of AI capabilities in the field.
keywords:
@@
journal: Astronomy &\& Computing
1 Introduction
Astronomy and its related fields demand sophisticated tools that can process vast amounts of specialized knowledge. Large Language Models (LLMs) have emerged as promising assistants for this domain, offering capabilities as research collaborators, educational resources, and knowledge repositories. Domain-specialized models demonstrate cost-effectiveness in such contexts, as their parameters can be optimized for specific knowledge domains rather than distributed across the breadth of general internet content. Our previous work, AstroSage-Llama-3.1-8B (de_haan_achieving_2025) (hereafter TdH25), established that an 8-billion parameter LLM, when trained on astronomical content, could match or exceed the performance of much larger general-purpose models on astronomical knowledge tasks (tingAstroMLab1Who2024). This finding highlighted the potential of domain specialization for creating efficient, high-performing AI assistants.
@@ 1 Introduction
This paper is structured as follows: Section 2 details the multi-stage training pipeline for AstroSage-Llama-3.1-70B, including improvements to our CPT and SFT corpora and procedures, and the model merging strategy for the final release. Section 3 discusses potential applications. Section 4 evaluates AstroSage-Llama-3.1-70B against the updated tingAstroMLab1Who2024 benchmark, comparing it with other recent proprietary and open-weight models.
2 Model Architecture and Training Methodology
AstroSage-Llama-3.1-70B is derived from the Meta-Llama-3.1-70B architecture (dubeyLlama3Herd2024b). This base model was selected for consistency with TdH25, which chose Meta-Llama-3.1-8B for its strong general capabilities and permissive licensing. The tokenizer from Meta-Llama-3.1-70B-Instruct was used without modification. Following the methodology of TdH25, our development process comprised three main stages: continued pre-training (CPT), supervised fine-tuning (SFT), and model merging.
2.1 Continued Pre-training (CPT)
The objective of CPT is to imbue the base model with domain-specific knowledge from astronomical literature. The CPT dataset for AstroSage-Llama-3.1-70B builds upon the comprehensive corpus developed in TdH25, which included approximately 250,000 arXiv preprints (astro-ph, gr-qc from 2007-2024), nearly 30,000 Wikipedia articles related to astronomy, and internet-available textbooks. The knowledge cutoff for the astro-ph papers included in the CPT dataset remains January 2024.
@@ 2.1 Continued Pre-training (CPT)
  • We applied ftfy (speer_ftfy_2019) for consistent Unicode text normalization and some rule-based repetition removal to correct OCR failures, supplementing the perplexity-based cleaning methods detailed in TdH25.
  • To preserve general language understanding and mitigate catastrophic forgetting due to specialization, we incorporated a random selection of samples from the FineWeb dataset (penedoRefinedWebDatasetFalcon2023) into each training epoch. This addition of previous pretraining tokens during CPT is sometimes known as “replay”. The specific FineWeb samples were varied for each epoch, ensuring diverse exposure to general web text.
The CPT phase was conducted on the Oak Ridge Leadership Computing Facility (OLCF) Frontier supercomputer utilizing AMD Instinct MI250X GPUs. The Meta-Llama-3.1-70B base model underwent 2.5 epochs of CPT, a training process that utilized approximately 168,000 GPU-hours. Several learning hyperparameters were modified compared to TdH25. Firstly, the learning rate was scaled relative to that in dubeyLlama3Herd2024b to maintain the same learning rate per token. Secondly, the warm-up period was extended to 0.15 epochs. Finally, we increased the weight decay regularization term to 0.1, aligning with dubeyLlama3Herd2024b.
2.2 Supervised Fine-tuning (SFT)
@@
Figure 1: Composition of the AstroSage-Llama-3.1-70B SFT training dataset. The combination of reasoning-focused datasets (22.7%) with domain-specific astronomy Q&A (22.0%) reflects our strategy to develop a model that combines analytical thinking with specialized knowledge. General instruction-following data (OpenHermes 2.5) helps maintain versatility while preserving domain expertise. File sizes represent uncompressed UTF-8 encoded text including system and user prompts.
Figure 1 illustrates the composition of the SFT dataset. Its largest component is NVIDIA’s Llama-Nemotron-Post-Training-Dataset (bercovich2025llamanemotronefficientreasoningmodels), which was used to train the Llama-3_1-Nemotron series. This series demonstrates strong performance on the LMArena (chiangChatbotArenaOpen2024), suggesting it is a strong dataset for eliciting reasoning and aligning with human preferences. The reasoning components cover science, code, mathematics, and general chat, providing a foundation for analytical thinking across different domains.
@@
To enhance domain expertise, we incorporated the custom domain-specific Q&A datasets from both TdH25 and de_haan_cosmosage_2025, which together comprise approximately 30% of the training data.
@@
The model was fine-tuned on this SFT dataset for 2 epochs. This phase consumed approximately 43,000 GPU-hours on the same MI250X GPU architecture. Hyperparameters mirrored those of the CPT stage, with the exception of weight decay, which was removed.
2.3 Training Implementation
@@
The CPT and SFT stages were conducted on the Oak Ridge Leadership Computing Facility (OLCF) Frontier supercomputer using AMD Instinct MI250X GPUs. Our implementation employed the GPT-NeoX framework (gpt_neox_library), which we adapted for compatibility with the Llama-3.1 architecture. Training was distributed across 2,048 Graphics Compute Dies (GCDs) using a multi-dimensional parallelism strategy: tensor parallelism 8, pipeline parallelism 8, and data parallelism 32. As GPT-NeoX does not currently support DeepSpeed ZeRO stage 2/3 with pipeline parallelism, we used ZeRO stage 1 (rasleyDeepSpeedSystemOptimizations2020) with activation checkpointing enabled. This configuration achieved a computational throughput of approximately 50 TFLOPS/s per GCD, consistent with performance metrics reported by dashOptimizingDistributedTraining2023.
@@
Figure 2: Training dynamics for continued pre-training (CPT) and supervised fine-tuning (SFT). The top panel shows the training loss trajectory across 3 epochs of CPT followed by 2 epochs of SFT. Despite an early spike during CPT Epoch 1, the loss steadily decreases throughout CPT with minimal discontinuity at epoch boundaries, indicating effective learning without overfitting. The SFT phase exhibits a rapid initial decrease in loss, and shows discontinuous behavior at the epoch boundary, indicating the different statistical properties of the SFT data. The bottom panel shows the learning rate schedule, including the initial warm-up period, several manual adjustments during early CPT, the planned cosine decay during the final partial CPT epoch, and a separate warm-up and decay cycle for the SFT phase.
@@
Figure 2 illustrates the training dynamics during both CPT and SFT phases. Several patterns emerge from these curves. The CPT loss exhibits a consistent downward trajectory without discontinuities at epoch boundaries—a positive indicator that the model was still learning effectively from the dataset without overfitting. The SFT loss shows a rapid initial decrease and exhibits discontinuous behavior at the epoch boundary, reflecting the different statistical properties of the SFT data compared to CPT.
@@
The SFT training loss initially exceeds the final CPT loss but decreases, dropping below 0.6. This indicates that, on average, the model assigns more than half its probability ($\left<P\right>=\exp(-\mathrm{loss})$) to the correct token—likely reflecting the more predictable nature of assistant responses compared to the CPT corpus. The model obtained after SFT is named AstroSage-70B-SFT.
2.4 Model Merging
@@
To create the final, publicly released AstroSage-Llama-3.1-70B, we employed model merging (also known as parameter averaging) using the mergekit library (goddardArceesMergeKitToolkit2024). This technique allows us to combine the strengths of our specialized SFT model with the robust instruction-following capabilities of other popular fine-tuned Llama-3.1-70B variants.
@@
The final mixture for AstroSage-Llama-3.1-70B was created using the DARE-TIES method (yuLanguageModelsAre2024). The AstroSage-70B-CPT model (our domain-specialized, pre-instruction-tuning model) was designated as the “base model” for the merge operation. The components and their contributions are:
  • 85% AstroSage-70B-SFT (density 0.9)
  • 15% Meta-Llama-3.1-70B-Instruct (density 0.5)
The final model inherits the astronomical expertise from AstroSage-70B-SFT while incorporating the general instruction-following and conversational polish from Meta’s Instruct model.
3 Features and Reasoning Capability
@@
A key feature of AstroSage-Llama-3.1-70B is its explicit reasoning capability. This aligns with recent advances in the broader LLM field, where explicit reasoning has emerged as a key development for enhancing model performance on complex tasks. The integration of reasoning mechanisms has become common in leading models, including OpenAI’s o1 through o4 series, Anthropic’s Claude models with “thinking mode,” DeepSeek-R1, and others. These developments demonstrate that exposing and structuring the internal reasoning process allows models to tackle complex problems systematically, resulting in improved accuracy and reliability.
@@
Building on these industry-wide insights, AstroSage-Llama-3.1-70B implements explicit reasoning that can be activated at inference time by setting the system prompt to detailed thinking on and prefilling the assistant completion with <think>. When enabled, the model generates a step-by-step reasoning process before providing the final answer to the user’s query. This is beneficial for complex astronomical problems requiring multi-step analysis. As the reasoning tokens are enclosed within <think> and </think> tags, they can easily be hidden from the end-user if desired.
4 Evaluation
4.1 Benchmark Dataset and Methodology
To evaluate AstroSage-Llama-3.1-70B, we utilize the AstroMLab-1 benchmark developed in tingAstroMLab1Who2024. Unlike the original study, which used an internal version, we evaluate against the officially released version hosted on Hugging Face (https://huggingface.co/datasets/AstroMLab/Astrobench_MCQ_v1_Public). The original dataset comprises 4,425 multiple-choice questions derived from Annual Review of Astronomy and Astrophysics (ARAA) papers that were explicitly withheld from the AstroSage training corpus.
To ensure all questions are answerable based on general astronomical knowledge, we applied a cleaning procedure using Claude-3.5-Sonnet. Questions requiring access to specific paper content (e.g., particular numerical results or methodology choices unique to that study) were removed, and questions with unnecessary paper references were revised. This reduced the evaluation set to 3,846 validated questions.
For testing, we provided all models with minimal system prompts indicating they should act as astronomy experts. For models with explicit reasoning modes (e.g., o1, DeepSeek-R1, Gemini with “thinking”), we enabled this feature. Reasoning modes did not improve scores for most models, likely because AstroMLab-1 questions primarily test knowledge recall rather than multi-step reasoning.
4.2 Model Performance
This evaluation significantly expands upon the initial benchmarking by tingAstroMLab1Who2024. Using the refined, validated subset of AstroMLab-1, we assess the performance of 119 state-of-the-art models, superseding the preliminary results published in July 2024.
To quantify cost-efficiency, we define a value score relative to the GPT-4 baseline (value = 0):
$\mathrm{Value}=\frac{\mathrm{Score}-\mathrm{Score}_{\mathrm{GPT\textrm{-}4}}}{3.5}+\log_{10}\left(\frac{\mathrm{Cost}_{\mathrm{GPT\textrm{-}4}}}{\mathrm{Cost}}\right)$ (1)
Here, the denominator 3.5 represents the empirical performance slope observed in tingAstroMLab1Who2024, where a tenfold increase in cost typically yields a 3.5 percentage point gain in accuracy. Intuitively, a Value Score of +1+1 signifies that a model delivers equivalent performance to the baseline while reducing inference costs by an order of magnitude ($10\times$). A score of +5+5 represents a $100,000\times$ improvement in cost-efficiency; for context, Gemini 3.0 Flash achieves a value of 4.98.
Table 1 presents the complete benchmark results organized by model family, reporting the score, model type (open-weight or proprietary), and value score.
Table 1: Benchmark results on the AstroMLab-1 dataset (3,846 questions). Value scores are computed relative to GPT-4 (value = 0). Type: O = Open-weight, P = Proprietary.
Model Score T Value Model Score T Value Model Score T Value
AstroSage (Astronomy-Specialized)
AstroSage-Llama-3.1-8B 81.9 O +4.12 AstroSage-Llama-3.1-70B 89.0 O +5.55
OpenAI/GPT Alibaba/Qwen Zhipu/GLM
GPT-3.5 72.7 P +0.21 Qwen 7B 59.2 O -2.12 GLM-3 Turbo 66.8 P -0.64
GPT-4 77.8 P 0.00 Qwen 1.5 7B 65.9 O -0.22 GLM-4 Flash 69.5 P +1.14
GPT-4o-mini 78.9 P +2.39 Qwen 1.5 14B 70.1 O +0.78 GLM-4 AirX 74.9 P +0.68
GPT-4o 85.1 P +2.94 Qwen 1.5 32B 75.4 O +1.98 GLM-4 Air 75.2 P +1.79
GPT-4.1-nano 80.8 P +3.11 Qwen 1.5 110B 75.4 O +1.50 GLM-4 0520 77.6 P +0.47
GPT-4.1-mini 86.7 P +4.19 Qwen 2 7B 69.9 O +0.92 GLM-4 Plus 80.5 P +2.43
GPT-4.1 86.6 P +3.47 Qwen 2 57B 73.9 O +1.26 GLM-4 9B 72.3 O +1.40
GPT-5-nano 83.1 P +3.82 Qwen 2 72B 80.5 O +3.03 GLM-4 32B 75.6 O +1.96
GPT-5-mini 86.7 P +4.15 Qwen 2.5 7B 72.5 O +1.31 GLM-4.5 Air 82.2 P +3.22
GPT-5 88.3 P +3.92 Qwen 2.5 14B 73.1 O +1.48 GLM-4.5 85.3 O +3.62
GPT-5.2 89.4 P +4.09 Qwen 2.5 32B 73.3 O +1.25 GLM-4.6 85.1 O +3.46
o1-mini 82.6 P +2.17 Qwen 2.5 72B 81.0 O +3.00 GLM-4.7 84.9 O +3.54
o1 86.2 P +2.49 Qwen 2.5 Max 83.5 P +2.69 Google/Gemini
o3-mini 86.2 P +3.64 Qwen 3 14B 79.0 O +2.89 Gemini 1.0 Pro 73.3 P +0.37
o3 87.2 P +3.64 Qwen 3 30B-A3B 81.2 O +3.50 Gemini 1.5 Flash 78.8 P +2.04
o4-mini 86.2 P +3.63 Qwen 3 32B 82.1 O +3.69 Gemini 1.5 Pro 80.5 P +1.59
gpt-oss-20b 84.0 O +4.67 Qwen 3 80B 83.2 O +3.85 Gemini 2.0 Flash 83.2 P +3.82
gpt-oss-120b 85.7 O +4.83 Qwen 3 235B 86.3 O +4.53 Gemini 2.0 Pro 85.9 P +3.26
Meta/Llama QwQ 32B 81.1 O +3.18 Gemini 2.5 Flash 83.9 P +3.27
Llama 2 7B 51.1 O -4.45 Anthropic/Claude Gemini 2.5 Pro 86.9 P +3.52
Llama 2 70B 72.6 O +0.89 Claude 2 77.8 P +0.47 Gemini 3.0 Flash 90.3 P +4.98
Llama 3 8B 75.2 O +2.22 Claude 3 Haiku 80.6 P +2.59 Gemini 3.0 Pro 88.9 P +3.99
Llama 3 70B 82.8 O +3.78 Claude 3 Sonnet 79.8 P +1.29 Google/Gemma
Llama 3.1 8B 75.7 O +2.36 Claude 3 Opus 85.0 P +2.07 Gemma 2B 44.9 O -6.05
Llama 3.1 70B 82.5 O +3.60 Claude 3.5 Haiku 78.0 P +1.33 Gemma 7B 57.5 O -2.60
Llama 3.1 405B 85.8 O +3.54 Claude 3.5 Sonnet 86.4 P +2.88 Gemma 2 2B 60.5 O -1.60
Llama 3.2 1B 47.3 O -5.34 Claude 3.7 Sonnet 86.3 P +3.13 Gemma 2 9B 73.7 O +1.72
Llama 3.2 11B 76.7 O +2.50 Claude 4 Sonnet 86.9 P +3.30 Gemma 2 27B 77.5 O +1.75
Llama 3.2 90B 82.9 O +3.51 Claude 4 Opus 87.8 P +2.87 Gemma 3 4B 69.1 O +0.56
Llama 3.3 70B 82.6 O +3.72 Claude 4.5 Haiku 83.3 P +2.75 Gemma 3 12B 76.2 O +2.39
Llama 4 Scout 83.6 O +4.13 Claude 4.5 Sonnet 87.3 P +3.42 Gemma 3 27B 78.7 O +2.91
Llama 4 Maverick 85.5 O +4.38 Claude 4.5 Opus 89.0 P +3.68 Microsoft/Phi
DeepSeek Mistral AI Phi-2 66.3 O +0.08
DeepSeek 67B 65.3 O -0.61 Mistral 7B 66.1 O -0.28 Phi-3 mini 73.6 O +1.99
DeepSeek V2 75.9 O +1.80 Mixtral 8x7B 76.7 O +1.61 Phi-3 medium 77.7 O +2.74
DeepSeek V2.5 78.7 O +2.77 Mixtral 8x22B 80.4 O +1.79 Phi-3.5 mini 74.6 O +2.28
DeepSeek V3 85.3 O +3.92 Mistral Small 3.1 81.3 O +3.82 Phi-4 mini 71.4 O +1.36
DeepSeek V3.1 85.8 O +4.30 Mistral Medium 3.1 84.0 P +3.35 Phi-4 82.2 O +3.91
DeepSeek V3.2 84.9 O +4.17 Mistral Large 1 79.3 P +2.09 xAI/Grok
DeepSeek R1 88.2 O +4.42 Mistral Large 2 83.1 P +2.57 Grok Beta 82.0 P +1.91
Moonshot/Kimi Mistral Large 3 85.0 O +3.73 Grok 2 84.0 P +2.67
Moonshot v1 74.6 P +0.52 MiniMax Grok 3 Mini 85.6 P +4.30
Kimi K2 87.5 O +4.23 MiniMax-01 83.4 P +2.55 Grok 4 Fast 85.6 P +4.35
MiniMax-M2 84.5 P +3.69
MiniMax-M2.1 83.8 O +3.90